# Frame-Voyager: Learning to Query Frames for Video Large Language Models

### v2

@@ 5 Conclusions
For the data construction, due to resource constraints, 1) we only generate the ranking data using VILA-8B, precluding experiments with more powerful Video-LLMs that might yield superior results. 2) The combinations for data construction are limited in size as the training set primarily focuses on short videos. Despite the promising improvements achieved, we highlight that applying Frame-Voyager to longer videos with a greater number of frame combinations could potentially lead to further performance enhancements in processing extended video content.
For the model framework, 1) we integrate our approach into existing Video-LLMs as a plug-in to ensure parameter efficiency, without additional fine-tuning of the backbone model. However, directly reusing the backbone’s parameters may not yield optimal results. Moreover, our simplified plug-in module could benefit from a more sophisticated design to further enhance overall performance. 2) Integrating our approach into the pre-training process of Video-LLMs remains unexplored, we hypothesize that learning frame combinations during pre-training could produce a more robust and effective model.
@@ Appendix A Implementation Details of Frame Extraction Baselines in RQ1
For the cluster-based approach, Katna, we directly use Katna API. It initially segments the videos according to content specifics and extracts a list of keyframes from each individual segment, employing the Histogram with K-means technique. Afterward, all segmented keyframe lists are merged to pick out the final $T$ frames.
For VILA-Embedding, we utilize the visual feature after the projector and the text feature from word embedding to measure the cosine similarity after the average pooling. The cosine similarity is then used for ranking the frames. For CLIP, we use the hosted model on HuggingFace [^3] https://huggingface.co/docs/transformers/model_doc/clip and directly rank and select top-$T$ candidate frames according to their output logits.
@@ Appendix B Question type analysis of the generated datasets
Appendix C Computation and Parameter Cost
@@
For inference, we randomly sample $100$ examples from Video-MME to measure the model’s inference latency. Using the experimental setting from our main paper, the average inference latency for the uniform sampling baseline is $1.329$ seconds per example, while for Frame-Voyager it is $1.696$ seconds, indicating that our method introduces an additional latency overhead of approximately $27.6$%.
Appendix D A Study on the Number of Candidate Frames
@@
As the number of candidate frames increases, more video information is captured within the selection pool. While selecting $8$ frames from a larger candidate set, our method consistently improves results on the Video-MME dataset. As shown in Table D, selecting $8$ frames from a candidate pool of $128$ frames yields a $3$% improvement compared to selecting from just $8$ frames (i.e., uniform sampling). However, as the candidate set size increases further (from $128$ to $256$), the additional information brought by the expanded set becomes limited.
Appendix E More Case Studies
@@
In Figure 9 and 10, we present additional case studies to highlight the effectiveness of our model. In Figure 9, the uniform sampling fails to capture key objects mentioned in the query. Similarly, the CLIP-based method struggles due to its limited OCR capabilities on special fonts, often match objects like “blue food dye”. In contrast, our Frame-Voyager, accurately identifies the used ingredients in this video, providing sufficient video context to answer the query about which ingredients are not used.

### v3

@@ 5 Conclusions
For the data construction, due to resource constraints, 1) we only generate the ranking data using VILA-8B, precluding experiments with more powerful Video-LLMs that might yield superior results. 2) The combinations for data construction are limited in size as the training set primarily focuses on short videos. Despite the promising improvements achieved, we highlight that applying Frame-Voyager to longer videos with a greater number of frame combinations could potentially lead to further performance enhancements in processing extended video content. For the model framework, 1) we integrate our approach into existing Video-LLMs as a plug-in to ensure parameter efficiency, without additional fine-tuning of the backbone model. However, directly reusing the backbone’s parameters may not yield optimal results. Moreover, our simplified plug-in module could benefit from a more sophisticated design to further enhance overall performance. 2) Integrating our approach into the pre-training process of Video-LLMs remains unexplored, we hypothesize that learning frame combinations during pre-training could produce a more robust and effective model.
@@ Appendix A Implementation Details of Frame Extraction Baselines in RQ1
For the cluster-based approach, Katna, we directly use Katna API. It initially segments the videos according to content specifics and extracts a list of keyframes from each individual segment, employing the Histogram with K-means technique. Afterward, all segmented keyframe lists are merged to pick out the final $T$ frames. For MDF (Han et al., 2024), we implement it utilizing the official code with default parameters [^3] https://github.com/declare-lab/Sealing. To compute the similarities between frames, we utilize the features from VILA-8B’s visual encoder.
For TempGQA (Xiao et al., 2024), we select a specific grounding segment based on the question following the official code [^4] https://github.com/doc-doc/NExT-GQA/tree/main/code/TempGQA. Then we uniformly sample frames from the selected segment and feed them into VILA-8B with questions to generate answers.
For SeViLA (Yu et al., 2024), we utilize its frame selection module, i.e., the localizer in SeViLA, and maintain the original hyperparameter settings [^5] https://github.com/Yui010206/SeViLA?tab=readme-ov-file.
For VILA-Embedding, we utilize the visual feature after the projector and the text feature from word embedding to measure the cosine similarity after the average pooling. The cosine similarity is then used for ranking the frames. For CLIP [^6] https://huggingface.co/openai/clip-vit-large-patch14-336 (Radford et al., 2021), SigLIP [^7] https://huggingface.co/google/siglip-so400m-patch14-384 (Zhai et al., 2023), and InternViT-6B (Chen et al., 2024), we directly rank and select top-$T$ candidate frames according to their output logits.
@@ Appendix B Question type analysis of the generated datasets
Appendix C Analysis of Computation Cost
@@
For inference, we randomly sample $100$ examples from Video-MME to measure the model’s inference latency. Using the experimental setting from our main paper, the average inference latency for the uniform sampling baseline is $1.329$ seconds per example, while for Frame-Voyager it is $1.696$ seconds, indicating that our method introduces an additional latency overhead of approximately $27.6$%. Meanwhile, Frame-Voyager uses $23,069$ MiB and the baseline uses $22,425$ MiB, making a difference of 2.9%.
Appendix D Results on More Benchmarks
In this section, we present additional results on three additional benchmarks, i.e., MVBench (Li et al., 2024d), STAR (Wu et al., 2021), and EgoSchema (Mangalam et al., 2023), in Table 4. The results show that our method consistently outperforms uniform sampling on the first two benchmarks, while the improvement of Frame-Voyager on EgoSchema is marginal. After carefully examining the EgoSchema dataset, we identify the main reasons are ambiguous pronouns and camera wearer (cameraman) related questions. Upon 50 random sampled instances from EgoSchema, we obverse that 44 instances contain the character “c” rather than the conventional pronoun to represent the human. For example, “what was the primary tool used by c in the video” and “what can be inferred about c’s assessment of the plants during the video”. We find that “c” represents the “camera wearer”, who may not even appear in the video. Such special characteristics impede Frame-Voyager’s ability to identify truly query-relevant frame combinations.
[…]
MVBench STAR EgoSchema
VILA-8B 40.1 48.3 53.3
+Frame-Voyager 41.1 50.5 53.6
VILA-40B 56.2 62.1 63.2
+Frame-Voyager 57.3 63.8 63.3
Appendix E A Study on the Number of Candidate Frames
@@
As the number of candidate frames increases, more video information is captured within the selection pool. While selecting $8$ frames from a larger candidate set, our method consistently improves results on the Video-MME dataset. As shown in Table E, selecting $8$ frames from a candidate pool of $128$ frames yields a $3$% improvement compared to selecting from just $8$ frames (i.e., uniform sampling). However, as the candidate set size increases further (from $128$ to $256$), the additional information brought by the expanded set becomes limited.
Appendix F Dynamic Frame Selection
In this section, we further explore the capability of Frame-Voyager on dynamic frame selection. Specifically, we modify the inference process by normalizing the rewards of candidate frames ranging from 00 to $1$. Then we only select the frames that exceed a specified threshold. We conduct experiments on Video-MME (without subtitles) using VILA-8B and LLaVA-OneVision-7B. For VILA-8B, we set a normalized reward threshold of $0.7$ and maintain our original constraints (maximum $8$ frames, minimum $1$ frame). For LLaVA-OneVision-7B, we expand the maximum frame limit to $32$ and adjust the threshold to $0.5$. The overall results is shown in Table 6.
Building on the experiment with VILA-8B, we further examine the required information density (defined by the number of dynamically selected frames) from three perspectives: video duration, video content domains, and query types. We present our experimental results below in Table 7, Table 8 and Table 9. The results reveal several patterns about how Frame-Voyager selection adapts to different scenarios:
  • Video Duration: Longer videos naturally require more frames (Short: 3.72, Medium: 4.07, Long: 4.61 frames on average).
  • Video Content Domains: The correlation between frame requirements and video domains is less obvious. Sports videos require the most average frames (4.43) due to frequent action changes, while other domains maintain relatively consistent frame requirements (3.89-4.31).
  • Query Types: Task complexity does influence frame selection. Complex tasks like counting (5.36 frames) and reasoning (4.58 frames) require more frames for comprehensive analysis. Note that counting tasks, despite having lower reasoning demands, usually requires answering questions like “How many different people appear in the video?”, which strongly depend on multiple frames. Simpler tasks like perception (3.01 frames) and recognition (3.84 frames) need fewer frames.
The results confirm Frame-Voyager’s effectiveness in dynamically selecting frames by considering both the video’s information density and the specific query needs.
[…]
Video-LLM Backbone Method Avg. Frames Acc.
VVILA-8B Uniform Sampling 4 44.7
Frame-Voyager + Fixed Num. 4 47.9
Uniform Sampling 8 47.5
Frame-Voyager + Fixed Num. 8 50.5
Frame-Voyager + Dynamic Num. 4.1 49.6
[…]
Frame-Voyager + Fixed Num. 8 57.5
Uniform Sampling 16 57.2
Frame-Voyager + Fixed Num. 16 59.2
Uniform Sampling 32 58.2
Frame-Voyager + Fixed Num. 32 60.4
Frame-Voyager + Dynamic Num. 13.3 59.1
Table 7: Averaged numbers of frames for different lengths of videos. The analysis is conducted based on the results of VILA-8B in Table 6.
Video Duration Avg. Frames
Short 3.72
Medium 4.07
Long 4.61
Table 8: Averaged numbers of frames for different video content domains. The analysis is conducted based on the results of VILA-8B in Table 6.
Video Content Domain Avg. Frames
Knowledge 3.89
Film & Television 3.96
Sports Competition 4.43
Artistic Performance 3.95
Life Record 3.97
Multilingual 4.31
Table 9: Averaged numbers of frames for different types of questions. The analysis is conducted based on the results of VILA-8B in Table 6.
Video Duration Avg. Frames
Perception 3.01
Recognition 3.84
Reasoning 4.58
OCR 3.93
Counting 5.36
Information Synopsis 4.35
Appendix G Optimization of Data Collection Pipeline
We implement two straightforward approaches to address efficiency concerns and reduce computational resource consumption in the data collection pipeline.
Method (1). We conduct dynamic pruning by filtering combinations based on frame-to-frame similarity and temporal proximities among frames. This strategy helps us discard the frame combinations containing multiple similar and redundant frames. It reduces the average number of combinations per sample to 14% on the VideoChatGPT dataset with processing time reduced to about 5 hours using 32 A100 GPUs and to 27% on the NextQA dataset with about 15 minutes using 8 A100 GPUs.
Method (2). We further utilize CLIP Radford et al. (2021) to compute the similarity between all frames and the question, and rank them accordingly. We mark lower-ranked frames as irrelevant. We believe that only a small portion of video frames directly relate to the question, while the majority of frames are irrelevant. As a result, most frame combinations only containing irrelevant frames can be discarded during the training process. Based on this pruning strategy, we filter out most combinations consisting of only irrelevant frames while retaining a few as low-ranking training samples. This approach can further reduce data collection time by 60-70%.
The comparison among the vanilla data collection method, +Method (1), and +Method (1) & (2) is shown below in Table 10. The backbone model is kept as VILA-8B. The results demonstrate that the data collection time of Frame-Voyager could be largely reduced to just 4.4% of the original version, while maintaining comparable performance across all benchmarks.
[…]
GPU Hours Time Costs Video-MME MLVU NextQA ANQA
Vanilla Method 1280+8 100% 50.5 49.8 51.1 51.6
+Method(1) 160+2 12.6% 50.6 50.0 50.8 52.0
+Method(1)&(2) 56+0.8 4.4% 50.5 49.7 51.0 51.9
Appendix H Case Study on Training Data
Figure 8: Case study from processed NextQA training data: These random samples show that lower loss values correlate with better visual-question alignment. The loss signal can enable learning to effectively select frame combinations for answering questions.
To evaluate the training objective of our Frame-Voyager, we examine random samples from the NextQA dataset after processing it through our data collection pipeline in Section 3.1. Figure 8 shows these sample cases, revealing a clear pattern: question-answer pairs with lower loss values show stronger alignment with the visual content. This correlation demonstrates that our loss-based ranking effectively identifies the most contextually relevant frames for answering questions. However, in cases with higher loss values, while the frames may contain relevant objects, they often lack sufficient visual information to fully answer the given questions.
Appendix I Case Study on Frame Combinations by Frame-Voyager
To address RQ6, we thoroughly examine the findings presented in Figure 9. In Figure 10 and 11, we present additional case studies to highlight the effectiveness of our model.
In Figure 10, the uniform sampling fails to capture key objects mentioned in the query. Similarly, the CLIP-based method struggles due to its limited OCR capabilities on special fonts, often match objects like “blue food dye”. In contrast, our Frame-Voyager, accurately identifies the used ingredients in this video, providing sufficient video context to answer the query about which ingredients are not used.